OpenStudy (anonymous):

4a=5-b and 4a=2-b Im confused! Please help!!

OpenStudy (sw050399):

Did you mean ln? As in the natural log?

OpenStudy (mathstudent55):

Is this a system of equations?

OpenStudy (mathstudent55):

4a = 5 - b 4a = 2 - b Subtract the second equation from the first equation: 0 = 3 + 0 0 = 3 Since 0 = 3 is a false statement, there is no solution.